OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x < -12 or x + 3 >= 12 A number line with an open circle on -6, shading to the left, and a closed circle on 9, shading to the right. A number line with a closed circle on -6, shading to the left, and an open circle on 9, shading to the right. A number line with an open circle on -6, a closed circle on 9, and shading in between. A number line with a closed circle on -6, an open circle on 9, and shading in between.

OpenStudy (anonymous):

@Notamathgenius

OpenStudy (anonymous):

still on this?

OpenStudy (notamathgenius):

Haleyy, I have no clue o_o sorry o_o

OpenStudy (anonymous):

yes i am .-.

OpenStudy (anonymous):

first is x<-6

OpenStudy (anonymous):

so open circle on -6 n shading on left

OpenStudy (anonymous):

2nd is x>=9

OpenStudy (anonymous):

so its closed circle n shading to right

OpenStudy (anonymous):

so what do itry to find as to be an answer?

OpenStudy (anonymous):

hmm consider 1st 1 2x<-12 so x< -12/2 x<-6 so all the terms less than -6 will be included xcept -6 

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? A) 3x - y = -7 B) 3x - y = -5 C) 3x - y = 5 D) x + 3y = -5

OpenStudy (anonymous):

and 2nd 1 is easy x+3>=12 x>=12-3 x>=9 so all terms greater than or equal to 12 

OpenStudy (anonymous):

but A was the answer?

OpenStudy (anonymous):

put values of (x,y) in equation

OpenStudy (anonymous):

A is the answer

OpenStudy (anonymous):

to the first one, but what about the second one?

OpenStudy (anonymous):

A put the values

OpenStudy (anonymous):

idk how.

OpenStudy (anonymous):

(-2,1) = (x,y) so put x= -2 and y=1

OpenStudy (anonymous):

3x - y = -7 in this

OpenStudy (anonymous):

if this doesnt satisfy then check for 2nd option

OpenStudy (anonymous):

you are supposed to check for (1, 10) too

OpenStudy (anonymous):

okay that just confused me.

OpenStudy (anonymous):

well then

OpenStudy (anonymous):

yupp